40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 9 12:35:47 2017, gautam, Update, ALS, Fiber ALS IMG_3942.JPG
    Reply  Mon Aug 14 16:24:09 2017, gautam, Update, ALS, Fiber ALS IMG_7471.JPGIMG_7472.JPGIMG_7473.JPGIMG_7474.JPG
       Reply  Wed Aug 16 20:24:23 2017, gautam, Update, ALS, Fiber ALS IMG_7476.JPGIMG_7477.JPGIMG_7478.JPGIMG_7479.JPG
          Reply  Wed Aug 23 17:22:36 2017, gautam, Update, ALS, Fiber ALS - reinstalled 
             Reply  Fri Aug 25 15:54:14 2017, gautam, Update, ALS, Fiber ALS noise measurement IMG_7497.JPGFOL_schematic.pdf20170825_IR_ALS.pdf
                Reply  Fri Aug 25 17:11:07 2017, rana, Update, ALS, Fiber ALS noise measurement 
                   Reply  Tue Sep 26 19:10:13 2017, gautam, Update, ALS, Fiber ALS setup neatened IMG_7605.JPG
                Reply  Sun Aug 27 11:57:31 2017, rana, Update, ALS, Fiber ALS noise measurement 
                   Reply  Tue Aug 29 02:08:39 2017, gautam, Update, ALS, Fiber ALS noise measurement 
                      Reply  Fri Sep 1 19:15:40 2017, gautam, Update, ALS, Fiber ALS noise measurement IR_ALS_noise.pdf
Message ID: 13246     Entry time: Wed Aug 23 17:22:36 2017     In reply to: 13222     Reply to this: 13254
Author: gautam 
Type: Update 
Category: ALS 
Subject: Fiber ALS - reinstalled 

I completed the revamp of the box, and re-installed the box on the PSL table today. I think it would be ideal to install this on one of the electronic racks, perhaps 1X2 would be best. We would have to re-route the fibers from the PSL table to 1X2, but I think they have sufficient length, and this way, the whole arrangement is much cleaner.

Did a quick check to make sure I could see beat notes for both arms. I will now attempt to measure the ALS noise with this revamped box, to see if the improved power supply and grounding arrangement, as well as fiber cleaning, has had any effect.

Photos + power budget + plan of action for using this box to characterize the green PDH locking to follow. 

For quick reference: here is the AM/PM measurement done when we re-installed the repaired Innolight NPRO on the new X endtable.

ELOG V3.1.3-